Homework Problems for Handout Sheet 28 In Problems 1 to 3, determine the eigenvalues and the corresponding solutions for theplease help in any of these questions in DE
Answer #1

Date: 1 1 Page No. To y tiy=0 auxillary 910150 equation )=0 y is caseo if i=0 > m=0.0 general solution is ya = AxtB. . y 12UPH y10)= 0= Date: 1 All 1 Page No. GILAS S0 = B.O - A=0 Ba cosait B Cosat = 0 if B=0 then we get trivial solution cosato o -RUPA Date: 1 Page No. e Auxiliary y101=0, y1=0 equation m2 + i=0 case o if d=0 - general solution is y01: ARAB ( from above eа Date: «5 18 тоо . - - - еекс 50\uКоо Кs (40) yWE AREX ylos - Beda Arete – Baeza (from Ex. 0 using given conditions. yol = 0Date: H Page: esing given conditions: 01-0 =) A y(L) to 8.0 -0 a med bacorab=0 a B cosal 20 Hor non-trivial solution = A= (